OpenStudy (anonymous):

step by step how do you do this problem; 3y-7+y+27=6y+20-4y

OpenStudy (anonymous):

Collect the Like terms...

OpenStudy (anonymous):

but how ive tryed that an its still telling me im wrong

OpenStudy (anonymous):

3y-7+y+27=6y+20-4y 3y+y+4y-6y=20+7-27 4y+4y-6y=27-27 8y-6y=0 2y=0 2y/2=0/2 y=0/2 that was my answer

OpenStudy (anonymous):

thank you